# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ 5.4 Divergent Association Task
Experimental Setup. The third use case considers the divergent association task (DAT) (Chen & Ding, 2023). Building on the theory that creativity is related to the capability of generating more divergent ideas (Beaty et al., 2014), it requires participants to name unrelated words. Then, their semantic distance can represent an objective measure of divergent thinking (Olson et al., 2021). DAT is a useful case study for decoding strategies as it constrains the generation to different nouns (thus, assuming an optimal probability distribution, the tail due to smoothing should contain everything else) and requires generating terms that are as different as possible, which is quite the opposite to what typically happens in language modeling: an optimal strategy should exclude non-appropriate tokens but also not limit too much the space of possible tokens. We strictly follow the setup proposed by Chen & Ding (2023). More concretely, given the embeddings of $n$ words, the DAT score is the average cosine distance between each pair of embeddings (then scaled as a percentage). We use GloVe embeddings (Pennington et al., 2014) and ask the model to generate a list of $10$ nouns. We discard outputs without at least $7$ distinct nouns and compute the DAT score for all other outputs over their first $7$ nouns. For completeness, Appendix D.1 reports results obtained when considering all $10$ nouns. We repeat the experiment $100$ times for non-greedy strategies to mitigate the sampling stochasticity.
@@
Experimental Results. Figure 2 summarizes the DAT results. For Meta-Llama-3-8B-Instruct, DiffSampling-cut has the highest average score (even if lower than the greedy score), and generates only valid outputs. Instead, both DiffSampling-lb and DiffSampling-minp perform slightly better or almost identically to the top-$p$ and min-$p$, respectively. The results for the pre-trained version Meta-Llama-3-8B are quite different. DiffSampling-cut is still substantially better than the greedy strategy, and it produces only valid outputs. However, all other methods achieve higher DAT scores while producing significantly fewer valid outputs. Although very similar, both DiffSampling-minp and DiffSampling-lb outperform their min-$p$ and top-$p$ counterparts, yielding either slightly higher scores or a greater number of valid outputs.
@@ 5.6 Temperature Scaling
Finally, we experiment with different temperature values $\tau$, i.e., $0.6$, 11, $1.5$, 22, and $10$. As detailed above, to preserve the mathematical guarantees of our approach, we apply temperature after the DiffSampling truncation, while our baselines apply this before (see Appendix D.2 for a comparison between temperature before and after truncation). As shown by Figure 3, DiffSampling+temperature preserves the output quality, and relevant differences only occur with our two relaxations and pre-trained models. Instead, the output quality rapidly drops with higher temperatures for the min-$p$ (by far the best of our baselines at $\tau>1.$) and top-$p$ baselines. In particular, the non-significant loss in quality for DiffSampling-cut confirms that our truncation strategy only preserves correct tokens. At the same time, temperature scaling has an (overall positive) impact on diversity; we refer to Appendix C for a detailed analysis of how all our quality and diversity metrics change at different $\tau$.
Figure 3: Average quality scores across different temperature values for top-$p$, min-$p$, and our methods.
@@ 6 Discussion
However, selecting the most appropriate method and hyperparameters is not straightforward and requires a case-by-case analysis. DiffSampling-cut works best when the task requires precision: whenever a user might otherwise rely on a greedy decoding strategy or a very low temperature, it enhances diversity without compromising accuracy. DiffSampling-lb fosters output diversity by trading off some accuracy, especially at higher values of $p_{lb}$ and, thus, appears most appropriate for divergent solutions. DiffSampling-minp is more well-balanced. Both can be used in place of top-$p$ and min-$p$ to “correct” them and potentially improve their diversity with no additional overhead. Increasing the temperature has proven highly effective for fine-tuned models across all methods, whenever it is not strictly necessary to preserve the original distribution.
Our experiments were limited to relatively small LLMs and based on quantitative, automatic evaluation. Several of the adopted metrics exhibit significant limitations (e.g., Schluter, 2017), often failing to align with human judgments (Tevet & Berant, 2021). Moreover, abstract concepts such as originality and creativity remain inherently difficult to define with precision (Franceschelli & Musolesi, 2025). We plan to experiment with human evaluators to verify whether the quality and diversity that DiffSampling aims to provide are also perceived by potential users.
@@ 7 Conclusion
In this paper, we have presented DiffSampling, a novel family of decoding strategies based on the analysis of the next-token distribution. In particular, given the distribution sorted in descending order, we compute the forward difference approximation of its discrete derivative and use it to remove tokens after its minimum value (possibly together with relaxations to allow for more diversity). In this way, we avoid incorrect tokens under the learned distribution. We have experimented with four different tasks, and our method has consistently performed at least as well as similar strategies in terms of accuracy, while potentially improving diversity.
@@
Our research agenda includes investigating whether combining DiffSampling with complementary techniques, such as re-ranking or controllable generation, can lead to further improvements in output quality. We also plan to leverage additional properties of the underlying probability distribution, beyond token likelihoods, to guide generation toward desired characteristics such as coherence, novelty, or user-specific preferences. These directions open up promising opportunities for enhancing the adaptability of text generation systems in general-purpose and task-specific settings.
Appendix A Computational Infrastructure and Implementation Details

### v5

@@ 5.4 Divergent Association Task
Experimental Setup. The third use case considers the divergent association task (DAT) (Chen and Ding, 2023). Building on the theory that creativity is related to the capability of generating more divergent ideas (Beaty et al., 2014), it requires participants to name unrelated words. In particular, the task is the following:
Write 10 nouns in English that are as irrelevant from each other as possible, in all meanings and uses of the words. Please note that the words you write should have only single word, only nouns (e.g., things, objects, concepts), and no proper nouns (e.g., no specific people or places).
Then, their semantic distance can represent an objective measure of divergent thinking (Olson et al., 2021). DAT is a useful case study for decoding strategies as it constrains the generation to different nouns (thus, assuming an optimal probability distribution, the tail due to smoothing should contain everything else) and requires generating terms that are as different as possible, which is quite the opposite to what typically happens in language modeling: an optimal strategy should exclude non-appropriate tokens but also not limit too much the space of possible tokens. We strictly follow the setup proposed by Chen and Ding (2023). More concretely, given the embeddings of $n$ words, the DAT score is the average cosine distance between each pair of embeddings (then scaled as a percentage). We use GloVe embeddings (Pennington et al., 2014) and ask the model to generate a list of $10$ nouns. We discard outputs without at least $7$ distinct nouns and compute the DAT score for all other outputs over their first $7$ nouns. For completeness, Appendix G.1 reports results obtained when considering all $10$ nouns. We repeat the experiment $100$ times for non-greedy strategies to mitigate the sampling stochasticity.
@@
Experimental Results. Figure 2 summarizes the DAT results, reporting both diversity (DAT score) and quality (count of valid outputs) measures. For Meta-Llama-3-8B-Instruct, the greedy strategy produces a strong set of nouns, achieving a DAT score higher than the average across all sampling methods. However, DiffSampling-cut generates a better set in almost half of the generations, and always produces a valid set of nouns. Instead, the other sampling schemes can produce better scores only occasionally, while sometimes failing at providing a valid set of nouns, and both DiffSampling-lb and DiffSampling-minp perform slightly better or almost identically to the top-$p$ and min-$p$, respectively. The results for the pre-trained version Meta-Llama-3-8B are quite different, and the quality-diversity trade-off is more pronounced. DiffSampling-cut is substantially better than the greedy strategy, and it produces only valid outputs. However, all other methods achieve higher DAT scores while producing significantly fewer valid outputs. Although very similar, both DiffSampling-minp and DiffSampling-lb outperform their min-$p$ and top-$p$ counterparts, yielding either slightly higher scores or a greater number of valid outputs.
@@ 5.6 Temperature Scaling
Finally, we experiment with different temperature values $\tau$, i.e., $0.6$, 11, $1.5$, 22, and $10$. As detailed above, to preserve the mathematical guarantees of our approach, we apply temperature after the DiffSampling truncation, while our baselines apply this before. Due to the different nature of temperature scaling, this comparison is intended only to highlight the impact of temperature position, rather than to imply that our method is superior to the baseline. To allow for a fairer analysis, we also report quality scores for DiffSampling when applying temperature before truncation (see Appendix F.5 for a full comparison between temperature before and after truncation). As shown by Figure 3, DiffSampling+temperature preserves the output quality, and relevant differences only occur with our two relaxations and pre-trained models. Instead, the output quality rapidly drops with higher temperatures for the min-$p$ (by far the best of our baselines at $\tau>1$) and top-$p$ baselines. In particular, the non-significant loss in quality for DiffSampling-cut confirms that our truncation strategy only preserves correct tokens. At the same time, temperature scaling has an (overall positive) impact on diversity; we refer to Appendix F for a detailed analysis of how all our quality and diversity metrics change at different $\tau$.
Figure 3: Average quality scores across different temperature values for top-$p$, min-$p$, and our methods when applying temperature before (dashed line) and after (full line) the truncation.
@@ 6 Discussion
However, selecting the most appropriate method and hyperparameters is not straightforward and requires a case-by-case analysis. If a small validation set is available, the choice of which strategy and parameters can be made empirically. Otherwise, our experiments show that DiffSampling-cut works best when the task requires precision: whenever a user might otherwise rely on a greedy decoding strategy or a very low temperature, it enhances diversity without compromising accuracy. DiffSampling-lb fosters output diversity by trading off some accuracy, especially at higher values of $p_{lb}$ and, thus, appears most appropriate for divergent solutions. DiffSampling-minp is more well-balanced. Both can be used in place of top-$p$ and min-$p$ to “correct” them and potentially improve their diversity with no additional overhead. Increasing the temperature has proven highly effective for fine-tuned models across all methods, whenever it is not strictly necessary to preserve the original distribution.
Our experiments were limited to relatively small LLMs, although preliminary results suggest that the same findings hold for larger models as well (see Appendix E for a more detailed analysis), and based on quantitative, automatic evaluation. Several of the adopted metrics exhibit significant limitations (e.g., Schluter, 2017), often failing to align with human judgments (Tevet and Berant, 2021). Moreover, abstract concepts such as originality and creativity remain inherently difficult to define with precision (Franceschelli and Musolesi, 2025). We plan to experiment with human evaluators to verify whether the quality and diversity that DiffSampling aims to provide are also perceived by potential users.
@@ 7 Conclusion
In this paper, we have presented DiffSampling, a novel family of decoding strategies based on the analysis of the next-token distribution. In particular, given the distribution sorted in descending order, we compute the forward difference approximation of its discrete derivative and use it to remove tokens after its minimum value (possibly together with relaxations to allow for more diversity). In this way, we avoid incorrect tokens under the learned distribution. We have experimented with four different tasks, and our method has consistently performed at least as well as similar strategies in terms of accuracy, despite sampling from a larger set of tokens, which has a positive impact on diversity.
@@
Our research agenda includes investigating whether combining DiffSampling with complementary techniques, such as re-ranking or controllable generation, can lead to further improvements in output quality. We also plan to leverage additional properties of the underlying probability distribution (e.g., its entropy (Potraghloo et al., 2025)), beyond token likelihoods, to guide generation toward desired characteristics such as coherence, novelty, or user-specific preferences. These directions open up promising opportunities for enhancing the adaptability of text generation systems in general-purpose and task-specific settings.
Broader Impact Statement
While our decoding scheme should, in theory, not increase the risk of generating tokens outside the true support, it may still produce unsafe content in certain contexts if the learned distribution itself is unsafe (e.g., containing learned biases, inappropriate language, or misleading information). Thus, it is important to continue using safety filters and domain constraints. Finally, we also perform a small check to ensure that DiffSampling does not increase unsafe content rates in the WritingPrompts use case, as it is the most open-ended generation task. Using Llama-Guard-3-8B (Inan et al., 2023), we found that the probability of unsafe outputs generated by our methods is identical to that of the corresponding methods at a temperature of $1.0$ or lower. However, the application of temperature after the truncation dramatically reduces the rate of unsafe generated text, especially for DiffSampling-cut and DiffSampling-minp. We report our results in Table 5.
Method Pre-Trained Instructed
placeholder $\tau=0.0$ $\tau=0.6$ $\tau=1.0$ $\tau=1.5$ $\tau=2.0$ $\tau=10.$ $\tau=0.0$ $\tau=0.6$ $\tau=1.0$ $\tau=1.5$ $\tau=2.0$ $\tau=10.$
Baselines placeholder placeholder
Greedy $12_{\pm.02} - - - - - 0.0$$04_{\pm.01} - - - - -$
Top-$p$ - $12_{\pm.01} 0.0$$09_{\pm.01}$ $0.62_{\pm.01}$ $0.70_{\pm.01}$ $69_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$ $0.57_{\pm.01}$ $0.69_{\pm.01}$ $0.71_{\pm.01}$
$\eta$-sampling - $1_{\pm.01} 0.11±.0$$1_{\pm.01}$ $0.72_{\pm.01}$ $0.70_{\pm.01}$ $68_{\pm.01} - 0.0$$04_{\pm.01}$ $0.11_{\pm.01}$ $0.62_{\pm.01}$ $0.69_{\pm.01}$ $0.69_{\pm.01}$
Locally typical - $12_{\pm.01} 0.0$$09_{\pm.01}$ $0.66_{\pm.01}$ $0.69_{\pm.01}$ $68_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$ $0.58_{\pm.01}$ $0.69_{\pm.01}$ $0.70_{\pm.01}$
Min-$p$ - $0.11_{\pm.01}$ $10_{\pm.01} 0.0$$09_{\pm.01}$ $0.19_{\pm.01}$ $68_{\pm.01} - 0.0$$5_{\pm.01} 0.0$$4_{\pm.01} 0.0$$04_{\pm.01}$ $0.30_{\pm.01}$ $0.69_{\pm.01}$
Ours ($\tau$ before) placeholder placeholder
DiffS.-cut - $0.13_{\pm.01}$ $0.12_{\pm.01}$ $0.12_{\pm.01}$ $0.12_{\pm.01}$ $58_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$04_{\pm.01}$ $0.47_{\pm.01}$
DiffS.-lb - $0.12_{\pm.01}$ $0.10_{\pm.01}$ $0.71_{\pm.01}$ $0.70_{\pm.01}$ $68_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$ $0.64_{\pm.01}$ $0.69_{\pm.01}$ $0.69_{\pm.01}$
DiffS.-minp - $0.11_{\pm.01}$ $10_{\pm.01} 0.0$$09_{\pm.01}$ $0.22_{\pm.01}$ $68_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$ $0.30_{\pm.01}$ $0.69_{\pm.01}$
Ours ($\tau$ after) placeholder placeholder
DiffS.-cut - $0.12_{\pm.01}$ $0.12_{\pm.01}$ $0.12_{\pm.01}$ $0.12_{\pm.01}$ $12_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$
DiffS.-lb - $0.11_{\pm.01}$ $0.10_{\pm.01}$ $0.44_{\pm.01}$ $0.52_{\pm.01}$ $51_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$05_{\pm.01}$ $0.32_{\pm.01}$ $0.37_{\pm.01}$ $0.40_{\pm.01}$
DiffS.-minp - $0.11_{\pm.01}$ $10_{\pm.01} 0.0$$9_{\pm.01} 0.0$$9_{\pm.01} 0.0$$8_{\pm.01} - 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$4_{\pm.01} 0.0$$5_{\pm.01} 0.0$$04_{\pm.01}$
Table 5: Unsafe probability of WritingPrompts outputs for baselines and our methods at different temperatures according to Llama-Guard-3-8B The mean and the $95\%$ confidence interval for the full set of answers are reported.
Appendix A Formal Analysis
In the following, we aim to formally define the conditions under which our truncation strategy is safe, i.e., the conditions under which all tokens up to $\operatorname*{arg\,min}\Delta p(x_{t}^{[n]})$ have a positive probability under the true distribution.
According to Hewitt et al. (2022), we can define the true probability distribution as $P_{\star}(\cdot|x_{<i})=\{p_{\star}^{[1]},\ldots,p_{\star}^{[|\mathcal{V}|]}\}$ with $\sum_{i=1}^{|\mathcal{V}|}p_{\star}^{[i]}=1$ and $p_{\star}^{[i]}\geq p_{\star}^{[i+1]}\;\forall\;i\in[1,|\mathcal{V}|-1]$, and where exists a $D<|\mathcal{V}|$ such that $p_{\star}^{[i]}=0\;\forall\;i>D$ (with $D=|\mathcal{V}|$, any truncation strategy is safe). Let us define a smoothing distribution $Q(\cdot|x_{<i})=\{q^{[1]},\ldots,q^{[|\mathcal{V}|]}\}$ with $q^{[i]}\in(\frac{1-\delta}{|\mathcal{V}|},\frac{1+\delta}{|\mathcal{V}|})\;\forall\;i\in[1,|\mathcal{V}|]$ and $\delta$ constant between 0 and $1$. Then, the learned distribution $P_{\theta}(\cdot|x_{<i})=\{p^{[1]},\ldots,p^{[|\mathcal{V}|]}\}$ with $\sum_{i=1}^{|\mathcal{V}|}p^{[i]}=1$ and $p^{[i]}\geq p^{[i+1]}\;\forall\;i\in[1,|\mathcal{V}|-1]$ can be defined as the weighted sum between the true probability distribution and the smoothing distribution:
$P_{\theta}(\cdot|x_{<i})=\lambda_{<i}P_{\star}(\cdot|x_{<i})+(1-\lambda_{<i})Q(\cdot|x_{<i})$ (2)
where $\lambda_{<i}\in(0,1]$. However, according to Hewitt et al. (2022), we can assume that $\lambda_{<i}\geq\max(\overline{\lambda}_{<i},\overline{\lambda})$, where $\overline{\lambda}_{<i}=1-\frac{|\mathcal{V}|\alpha\exp^{-H_{<i}}}{1+\delta}$ with $H_{<i}$ entropy of $P_{\star}(\cdot|x_{<i})$, and $\overline{\lambda}$ constant close to $1$; for simplicity, we will follow Hewitt et al. (2022) and assume to have $\overline{\lambda}=0.8$. This has two implications. First, the contribution provided by the smoothing distribution is bounded by $\alpha\exp^{-H_{<i}}$ with $\alpha\in[0,1]$ and generally very small, so the actual contribution depends on the entropy of the true distribution. Second, the weighting factor $\lambda_{<i}$ has a lower bound equal to $\overline{\lambda}=0.8$.
In this article, we propose to truncate the learned probability distribution at an index $K$ such that $K=\operatorname*{arg\,min}_{i}(p^{[i+1]}-p^{[i]})=\operatorname*{arg\,max}_{i}(p^{[i]}-p^{[i+1]})$. The truncation is safe when $K\leq D$, i.e., if the truncation only preserves tokens with a non-zero probability under the true distribution.
Proposition 1.
Given a learned probability distribution $P_{\theta}(\cdot|x_{<i})=\{p^{[1]},\ldots,p^{[|\mathcal{V}|]}\}$ sorted in descending order, the truncation performed by means of $\operatorname*{arg\,max}_{i}(p^{[i]}-p^{[i+1]})$ only preserves tokens from the support of the true distribution $P_{\star}(\cdot|x_{<i})$ if $\max_{i}(p_{\star}^{[i]}-p_{\star}^{[i+1]})>\frac{1}{|\mathcal{V}|}$.
Proof.
The truncation is safe whenever $K\leq D$, i.e., whenever the maximum difference between a token from the true support and its next token is greater than the maximum difference between tokens from outside the true support. The maximum difference between tokens from outside the true support, i.e., with $p_{\star}^{[i]}=p_{\star}^{[i+1]}=0$, is:
$\begin{split}\max_{i>D}(p^{[i]}-p^{[i+1]})&=\max_{i>D}((1-\lambda_{<i})q^{[i]}-(1-\lambda_{<i})q^{[i+1]})\\ &=(1-\lambda_{<i})(\frac{1+\delta}{|\mathcal{V}|}-\frac{1-\delta}{|\mathcal{V}|})=(1-\lambda_{<i})\frac{2\delta}{|\mathcal{V}|}.\end{split}$ (3)
Instead, the maximum difference between a token from the true support and its next token is given by:
$\begin{split}\max_{i\leq D}(p^{[i]}-p^{[i+1]})&=\max_{i\leq D}(\lambda_{<i}p_{\star}^{[i]}+(1-\lambda_{<i})q^{[i]}-\lambda_{<i}p_{\star}^{[i+1]}-(1-\lambda_{<i})q^{[i+1]})\\ &=\max_{i\leq D}(\lambda_{<i}(p_{\star}^{[i]}-p_{\star}^{[i+1]})+(1-\lambda_{<i})(q^{[i]}-q^{[i+1]})).\end{split}$ (4)
This value is lower-bounded by $\max_{i\leq D}(\lambda_{<i}(p_{\star}^{[i]}-p_{\star}^{[i+1]}))+\min_{i\leq D}((1-\lambda_{<i})(q^{[i]}-q^{[i+1]}))=\lambda_{<i}\max_{i\leq D}(p_{\star}^{[i]}-p_{\star}^{[i+1]})+(1-\lambda_{<i})\min_{i\leq D}(q^{[i]}-q^{[i+1]})$. The second term is exactly the opposite of the maximum value computed above: $\min_{i}(q^{[i]}-q^{[i+1]})=\frac{1-\delta}{|\mathcal{V}|}-\frac{1+\delta}{|\mathcal{V}|}=-\frac{2\delta}{|\mathcal{V}|}$. If we define $\Delta_{\star}^{[i]}=\max_{i\leq D}(p_{\star}^{[i]}-p_{\star}^{[i+1]})$, we obtain a lower-bounded maximum given by $\lambda_{<i}\Delta_{\star}^{[i]}-(1-\lambda_{<i})\frac{2\delta}{|\mathcal{V}|}$.
To have $K\leq D$, we impose:
$\begin{split}\lambda_{<i}\Delta_{\star}^{[i]}-(1-\lambda_{<i})\frac{2\delta}{|\mathcal{V}|}&>(1-\lambda_{<i})\frac{2\delta}{|\mathcal{V}|}\\ \lambda_{<i}\Delta_{\star}^{[i]}&>(1-\lambda_{<i})\frac{4\delta}{|\mathcal{V}|}\\ \Delta_{\star}^{[i]}&>\frac{4\delta}{|\mathcal{V}|}\frac{(1-\lambda_{<i})}{\lambda_{<i}}\end{split}$ (5)
Since $\lambda_{<i}$ is lower-bounded by $\overline{\lambda}$, the second term can be reduced to $\frac{(1-\overline{\lambda})}{\overline{\lambda}}$. As suggested by Hewitt et al. (2022), we assumed $\overline{\lambda}=0.8$; we obtain that our truncation strategy is safe if $\Delta_{\star}^{[i]}>\frac{4\delta}{|\mathcal{V}|}\frac{1}{4}=\frac{\delta}{|\mathcal{V}|}$. Since $\delta$ is upper-bounded to $1$, we get a lower bound of $\Delta_{\star}^{[i]}>\frac{1}{|\mathcal{V}|}$ that proves our proposition. ∎
Proposition 2.
Given a learned probability distribution $P_{\theta}(\cdot|x_{<i})=\{p^{[1]},\ldots,p^{[|\mathcal{V}|]}\}$ sorted in descending order, the truncation performed by means of $\operatorname*{arg\,max}_{i}(p^{[i]}-p^{[i+1]})$ preserves tokens only from the support with size $D$ of the true distribution $P_{\star}(\cdot|x_{<i})$ if $D<\sqrt{2|\mathcal{V}|}$.
Proof.
The maximum difference $\max_{i}(p_{\star}^{[i]}-p_{\star}^{[i+1]})$ is lower-bounded by $\frac{2}{D(D+1)}$. This bound holds when the differences between the first $D+1$ tokens (with the first $D$ tokens having a positive probability and the $D+1$-th zero probability) are equal, i.e., when the first $D+1$ tokens are equidistant. According to Proposition 1, $\operatorname*{arg\,max}_{i}(p^{[i]}-p^{[i+1]})\leq D$ if $\max_{i}(p_{\star}^{[i]}-p_{\star}^{[i+1]})>\frac{1}{|\mathcal{V}|}$. Thus, this is also true when $\frac{2}{D(D+1)}>\frac{1}{|\mathcal{V}|}$, i.e., if $D(D+1)<2|\mathcal{V}|$, providing an upper bound for the true support size $D$ of $\approx\sqrt{2|\mathcal{V}|}$, which proves our proposition. ∎
In summary, our truncation strategy is safe whenever we have $\max(p_{\star}^{[i]}-p_{\star}^{[i+1]})>\frac{1}{|\mathcal{V}|}$ or $D<\sqrt{2|\mathcal{V}|}$. To provide a practical intuition of the meaning of these conditions, consider the case of an LLM, whose vocabulary size $|\mathcal{V}|$ is usually in the order of $50000$. This sets the lower bound for the maximum difference at $0.00002$ and the upper bound for the true support size at $316$. If either of these conditions is satisfied, our truncation strategy is considered safe.
However, it is important to note that these two bounds represent only worst-case scenarios. First of all, we assumed that $\delta$ can be equal to $1$ and $\lambda_{<i}=\overline{\lambda}=0.8$. In practice, $\delta$ will be much smaller than $1$, leading to a proportionally smaller lower bound for $\max(p_{\star}^{[i]}-p_{\star}^{[i+1]})$. Moreover, $\lambda_{<i}$ will also assume the value of $\overline{\lambda}_{<i}$, which is likely to be higher than $\overline{\lambda}$ when $H_{<i}$ is higher, i.e., for very homogeneous distributions (unbalanced distributions most likely have a $\max(p_{\star}^{[i]}-p_{\star}^{[i+1]})>\frac{1}{|\mathcal{V}|}$). These two elements would make a realistic bound for $\max_{i}(p_{\star}^{[i]}-p_{\star}^{[i+1]})$ much looser. In addition, the proof of Proposition 2 considers an extremely conservative (and impossible to find in practice) lower bound for $\max_{i\leq D}(p^{[i]}-p^{[i+1]})$. Indeed, if all probabilities are equidistant, $\max_{i\leq D}(p^{[i]}-p^{[i+1]})$ is not lower-bounded by $\lambda_{<i}\frac{2}{D(D+1)}-(1-\lambda_{<i})\frac{2\delta}{|\mathcal{V}|}$ but by $\lambda_{<i}\frac{2}{D(D+1)}$. While the latter does not provide a valid lower bound for $\max_{i\leq D}(p^{[i]}-p^{[i+1]})$, the fact that the used lower bound cannot actually be attained, combined with the considerations regarding $\delta$ and $\lambda_{<i}$ discussed above, makes the upper bound for $D$ much looser in practice.
Appendix B Computational Infrastructure and Implementation Details